OpenStudy (anonymous):

Part 1 − Find the vertex, axis of symmetry, domain, and range of the graph of y = −3x2 − 3x + 4. Show all work for full credit. Part 2 − Using complete sentences, explain how you can determine the axis of symmetry, the domain, and range without graphing y = −3x2 − 3x + 4.

OpenStudy (e.cociuba):

Part 1: Vertex: (-½, 19/4) Axis of symmetry: x = -½ Domain: All real values Range: y ≤ 19/4

OpenStudy (anonymous):

ohh, awesome!! do you know part 2? @e.cociuba

OpenStudy (e.cociuba):

1 and 2. Use x = -b/2a to find the axis of symmetry. Here, a = -3, b = -3 and c = 4. Then, x = -(-3)/(2*-3) = -½ Axis of symmetry: x = -½ Take that coordinate and substitute that for the function to find the vertex. y(-½) = -3(-½)² - 3(-½) + 4 = 4.75 Vertex: (-0.5, 4.75) Domain: All real values Range: {y | y ≤ 4.75} since a < 0 for y = ax² + bx + c.

OpenStudy (anonymous):

DOMAIN IS ALL REAL NUMBERS SINCE THERE IS NO VALUE OF X FOR WHICH F(X) IS NOT DEFINED

OpenStudy (anonymous):

im mad WAH!!!!

OpenStudy (e.cociuba):

Umm ok? @dinosour1480

OpenStudy (anonymous):

thanks!! this is totally awesome guys!!!

OpenStudy (e.cociuba):

Np :))

OpenStudy (anonymous):

Thanks @MotherOfGod and @e.cociuba

OpenStudy (anonymous):

WASUUUUP